OpenStudy (anonymous):

Find x if (2/7)^2x = (7/2)^3x + 5.

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

good luck. try a graphing calculator

OpenStudy (anonymous):

how to solve ?

OpenStudy (anonymous):

no idea. can't use algebra. graph and see where they intersect

OpenStudy (zarkon):

Should the problem be \[\left(\frac{2}{7}\right)^{2x}=\left(\frac{7}{2}\right)^{3x+5}\] ?